about the universe forum commander Shop Now Commanders Circle
Product List FAQs home Links Contact Us

Monday, March 15, 2010

This week at ADB, Inc., 7-13 March 2010

Steve Cole reports:

The weather was mild all week, in the 50s and 60s. We had some rain early in the week.

The spam storm continued most of the week, with the on-deck filters stopping up to ten thousand of them per day. Saturday, it suddenly stopped. At the office, the floors were cleaned and polished this week.

I spent most of this week doing after-action reports for the F&E 2010 project. I did get Joel everything he needed to send Hailing Frequencies on Wednesday and we got Communique uploaded. I was sick most of Monday. I set up the file for Communique #52 and started doing stuff for it, such as updating the Gorn BDD.

We had a lot of mail orders this week, which kept Leanna and Mike Sparks busy. Leanna had cataract surgery on Friday; it went well and her vision is much improved.

Steve Petrick worked on CL41, R12, and C3A.

Joel Shutts got the newsletter released, Communique uploaded, the pull-down menus fixed on the Commander's Circle, and some stuff uploaded.

The contractor working on our house finished Phase Two (remodeling the master bedroom) which involved replacing the carpet with ceramic tile, removing the popcorn, repainting, replacing all of the woodwork including the doors, turning Leanna's old bathroom into my new walk-in closet and gun room, and added a modern ceiling fan. Next week, he begins Phase Three, which will remodel the rest of the house.